one x-intercept for a parabola is at the point (1,0). use the factor method to find the other x-intercept for the parabola defined by this equation: y=4x^2+4x-8
Answer by ewatrrr(24785)


Hi,
factoring to find x when y = 0
4x^2 + 4x - 8 = 0
4[x^2 + x - 2} = 0
4[(x-1)(x+2)]= 0 Note:SUM of the inner product(-x) and the outer product(2x) = x
(x-1)(x+2)= 0
(x-1)= 0
x = 1 already determined: Pt(1,0) named an x-intercept
(x+2)= 0
x = -2 Pt(-2,0) the other x-intercept